OpenStudy (anonymous):

1.Is the sequence geometric? If so, identify the common ratio. 6, 12, 24, 48, ... yes; 2 yes, –2 yes; 4 no 2.What is the fifth term of the geometric sequence? 5, 15, 45, ... 1,215 1,875 405 3,645

OpenStudy (anonymous):

1.A 2.i think might be B

OpenStudy (anonymous):

1) yes it is geometric

OpenStudy (anonymous):

the common ration is 2

OpenStudy (anonymous):

a geometric sequence is a sequence of numbers. that have multiplication or division in common. easy way to find out is take a number. and divide it by the number before. 24/12 = 2. now do the same for another number 48/24= 2. if they are equal. then it is geometric.

OpenStudy (anonymous):

explanations are better than answers fyi

OpenStudy (anonymous):

15/5 = 3 45/15= 3 so 45*3 gets the next answer

OpenStudy (anonymous):

243

OpenStudy (anonymous):

the fifth term is 243

OpenStudy (anonymous):

i mean the next number. than that number times 3. and so on

OpenStudy (anonymous):

so its C. cause 135*3 = 405

OpenStudy (anonymous):